Solve by Substitution,
I need a help please!
7x-8y=-1
6x+36=y
7x-8(6x+36)=-1
7x-48x-288= -1
41x=287
x=-7
AND
y=???? How??
Thank you!

Found 2 solutions by Alan3354, RAY100:
Answer by Alan3354(69443)

need a help please!
7x-8y=-1
6x+36=y
7x-8(6x+36)=-1
7x-48x-288= -1
41x=287
x=-7
AND
y=???? How??
--------------
y = 6x+36
= -42 + 36
= -6
-----
You were almost done.

Answer by RAY100(1637)

Very good work,,,,,(+41 should be -41)
.
with x=-7,,,substitute back into 1st eqn
.
7(-7) -8y = -1
.
-49 -8y = -1
.
-8y = -1 +49 = =48
.
y = -6
.
(-7,-6) is answer
.
check,,,(1) 7(-7) -8(-6) = -1,,,-49 +48 =-1,,,,ok
.
(2) 6(-7) +36 = -6,,,,-42 + 36 = -6,,,,,ok
.
Good work,,,Keep it going